OpenStudy (anonymous):

Explain why the x-coordinates of the points where the graphs of the equations y = 4^x and y = 2^(x + 2) intersect are the solutions of the equation 4^x = 2^(x + 2)

OpenStudy (anonymous):

This is a good thinking problem. What does it mean when two lines intersect?

OpenStudy (anonymous):

Uh, I don't know how to answer that other than where two lines meet on a coordinate plain

OpenStudy (anonymous):

That's good enough, and what happens at the intersection that is unique?

OpenStudy (anonymous):

I don't know what you mean, but I suppose it's unique because its the only place they meet?

OpenStudy (anonymous):

It's unique because that is the only place where the coordinates of the first line and the coordinates of the second line are exactly the same

OpenStudy (anonymous):

So x1=x2, and y1=y2

OpenStudy (anonymous):

Now look at the example they gave you. We will label the first as y1 and the second as y2. y1=4^x y2=2^(x+2)

OpenStudy (anonymous):

If we set 4^x and 2^(x+2) equal to each other, then that must also mean that y1=y2 right?

OpenStudy (anonymous):

So there is the first half of the intersection, now x1 needs to equal x2.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Hmm I'm having a tough time explaining this part, gimme a sec

OpenStudy (anonymous):

Hah, okay. :)

OpenStudy (anonymous):

this is because its the only point for which y have the same value for both the function... for a particular value of x

OpenStudy (anonymous):

What..?

OpenStudy (anonymous):

Oh, is that the answer?

OpenStudy (anonymous):

Oh okay I got it now. You want to find the x value that makes both sides equal, since that automatically means that all x's in both sides are equal.

OpenStudy (anonymous):

its the only point ... for which you will get value for both the functions same.... for a given particular value of x... this particular value is called the solution of x in the equation

OpenStudy (anonymous):

So that will determine what y is, and when y1 is also equal to y2, 4^x=y1=2^(x+2)=y2

OpenStudy (anonymous):

The explanation is a bit lengthy, so I'll try to summarize it. Both of those equations are equal to some y. When those y's are the same along with the same x's, you have an intersection.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Imagine plugging in random x's into both equations. Whichever yields matching y's is the answer

OpenStudy (anonymous):

Okay, thank you.